Malone Imports stock should return 12 percent in a boom, 10 percent in a normal economy, and 2 percent in a recession. The probabilities of a boom, normal economy, and recession are 5 percent, 85 percent, and 10 percent, respectively. What is the variance of the returns on this stock? Please explain your answer.
Answer #1

Expected return=Respective return*Respective probability

=(0.05*12)+(0.85*10)+(0.1*2)=9.3%

probability Return probability*(Return-Expected Return)^2
0.05 12 0.05*(12-9.3)^2=0.3645
0.85 10 0.85*(10-9.3)^2=0.4165
0.1 2 0.1*(2-9.3)^2=5.329
Total=6.11%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=(6.11)^(1/2)

=2.47%(Approx).

Variance=Standard deviation^2

=6.11%